OpenStudy (anonymous):

show all work in dividing y^2-y-12/y^2-6y+8 by y^2+2y-15/y^2-7y+10, use complete sentences to explain how to simplify this expression.

OpenStudy (amistre64):

it would help to factor them; flip the underone, and cancel like terms

OpenStudy (anonymous):

Factor them (y-4)(y+3)/(y-2)(y-4) Divided by (y-3)(y+5)/(y-5)(y-2) then cancel common terms Multiply first (y-4)(y+3)/1x1/(y-2)(y-4) =y+3/y-2 (y-3)(y+5)/1x1/(y-5)(y-2) =

OpenStudy (anonymous):

u must continue it

OpenStudy (anonymous):

Thank You!!!

OpenStudy (anonymous):

your welcome